PSa 2-9 Calculate Gross Pay Based on an Incentive Plan

Calculate gross pay for each of the following employees of Launchpad Co. The company offers a regular wage rate of $8.20/hour to all employees. Under an incentive plan in place for all employees, this rate increases for any employee who can meet weekly production goals. The increased rates and corresponding thresholds that must be met are as follows:

  • $9.40/hour for producing at least 2,000 units
  • $10.60/hour for producing at least 2,800 units
  • $11.80/hour for producing at least 3,700 units
  • $13/hour for producing at least 4,700 units

All employees are paid an overtime wage rate that is 1.5 times their respective regular wage rates.

NOTE: For simplicity, all calculations throughout this exercise, both intermediate and final, should be rounded to two decimal places at each calculation.

1: Bronson Chau worked 45 hours and produced 3,251 units.

Gross Pay = $

2: Pauline Myers worked 48 hours and produced 2,054 units.

Gross Pay = $

3: Angela Smith worked 54 hours and produced 5,157 units.

Gross Pay = $

4: Angelo Balducci worked 48 hours and produced 3,986 units.

Gross Pay = $
Answer #1

As per Fair Labor Standards Act (FLSA), overtime pay starts if an employee works for more than 40 hours a week.

Overtime rate= 1.5 times their respective regular wage rate

1. Bronson chau's wages

Wage rate(3251 units)= $10.6 per hour

Overtime rate= 1.5*10.6=$15.9 per hour

Regular wages= 10.6*40=$421

Overtime wages=15.9*5=$79.5

Gross pay= $500.5

2. Pauline Myers(2054 units)

Wage rate= $9.4 per hour

Overtime rate= $14.1 per hour

Regular wages=9.4*40= $376

Overtime wages=14.1*8=$112.8

Gross pay=$488.8

3. Angela Smith (5157 units)

Wage rate= $13 per hour

Overtime rate=13*1.5=$19.5 per hour

Regular wages=13*40=$520

Overtime wages= 19.5*14=$273

Gross pay=$793

4. Angelo Balducci(3986 units)

Wage rate=$11.8 per hour

Overtime rate=11.8*1.5=$17.7 per hour

Regular wages=11.8*40= $472

Overtime wages=17.7*8=$141.6

Gross pay=$613.6

Note: overtime= hours worked-40 hours.

> The first one I got 503.50 because the 10.6*40 is actually 424

Bayli Jeffords Tue, Feb 8, 2022 8:25 PM